# ECMSim: A high-performance web simulation of cardiac ECM remodeling through integrated ODE-based signaling and diffusion

### v1

@@ Introduction
[width=1]Fig-4.pdf Screenshots of 3 distinguish cases of simulation. All the matrices have collagen I in the ECM and simulated 4 different iterations (0, 100, 300, 500) with 0.1 time steps (S0, simulation times 0, 10, 30, and 50 respectively). More yellow color represents the higher concentration and black color represent lowest concentration in the matrix. Case 01: No brush selected. So, the inputs molecules have default values (all at 0). Case 02: Brush selected (X shape). TGF- input value is 1 in the brushed area. Rest of the inputs are in default values (0). Unselected area has default input molecule concentration. Case 03: Brush selected (X shape). TGF-, AngII, IL6, IL1, TNF, and NE input molecules are set at 1. Unselected area has default input molecule concentration. Each brush selected cases (2 and 3 - each case separately) included same brush selection throughout the continuous temporal iterations.
@@ Results
ECMSim successfully implements a comprehensive computational framework for simulating cardiac ECM remodeling through the integration of over 125 molecular species distributed across a spatially-resolved tissue grid. Figure 2A illustrates the fundamental feedback loop architecture underlying the model, where external TGF- input triggers intracellular signaling cascades that ultimately produce both ECM molecules and feedback molecules that diffuse to neighboring cells. This autocrine and the paracrine signaling mechanism enables the propagation and amplification of fibrotic responses through the tissue. The complete signaling network architecture (Figure 2B) encompasses the major pathways implicated in cardiac fibrosis, including TGF-/Smad signaling, MAPK cascades, PI3K-Akt-mTOR pathways, and mechanotransduction networks. The hierarchical organization of the network enables signal integration from multiple upstream inputs through intermediate processing nodes to downstream effector molecules responsible for matrix protein synthesis. The modular architecture facilitates systematic investigation of pathway crosstalk and the identification of critical regulatory nodes that control fibrotic responses.
The spatial implementation (Figure 2C) demonstrates the platform’s capability to model tissue-level organization through a regular cellular grid where each cell contains the complete intracellular signaling network. Intercellular communication is mediated through the diffusion of five key feedback molecules that coordinate regional responses and enable the emergence of spatially heterogeneous patterns. The diffusion dynamics (Figure 2D) show representative simulation outputs displaying the spatial distribution of molecular concentrations, illustrating how localized stimuli can propagate across the tissue domain through diffusive coupling.
@@
The ECMSim web application provides an intuitive interface for exploring complex signaling dynamics through multiple integrated visualization modalities (Figure 3). The primary heatmap visualization displays the spatial distribution of selected molecular species across the 100×100 cellular grid, with real-time updates during simulation execution that enable immediate observation of evolving concentration patterns. The molecule selection panel allows users to dynamically switch between 17 ECM molecules and 4 feedback molecules, instantly updating both spatial visualizations and temporal plots without requiring simulation restart. The innovative brush selection tool enables precise spatial control over input conditions by allowing users to define arbitrary cell populations for targeted interventions. The brush interface supports adjustable sizing from single-cell precision to broad regional selections, accommodating diverse experimental scenarios ranging from localized growth factor application to regional mechanical stimulation. The input concentration control panel provides independent adjustment of 10 primary signaling molecules, with real-time value displays and immediate propagation to selected cell populations. This design enables complex experimental scenarios such as spatially heterogeneous inflammation conditions or combinatorial cytokine treatments.
The dual-scale visualization system combines tissue-level spatial patterns with detailed single-cell temporal dynamics. Users can select up to two individual cells from the smaller overview grid, with their molecular trajectories displayed in real-time line plots that enable direct comparison between cells experiencing different microenvironments. The comprehensive parameter control panel contains all eight primary rate constants with real-time adjustment capabilities, enabling immediate exploration of model sensitivity and behavior under different kinetic assumptions. This interactive capability transforms ECMSim into a digital laboratory environment where hypotheses can be tested and validated without the temporal and resource constraints of traditional experimental approaches.
@@
The platform’s spatial modeling capabilities enable investigation of emergent tissue-level phenomena through the simulation of diffusion-driven pattern formation. Figure 4 demonstrates three representative simulation scenarios illustrate the platform’s ability to model spatially heterogeneous responses to localized stimuli. Case 01 represents the baseline condition with uniform zero input concentrations across all cells, resulting in spatially homogeneous concentrations dominated by stochastic fluctuations around the initial random values. The temporal evolution from iteration 0 to 500 shows gradual equilibration toward steady-state values with minimal spatial organization.
Case 02 demonstrates the response to localized TGF- stimulation applied in an X-shaped pattern through the brush selection interface. The TGF- input concentration of 1.0 in the selected region triggers robust intracellular signaling cascades that drive procollagen I (proCI) synthesis in the stimulated cells. The temporal evolution reveals the gradual emergence of spatial patterns that reflect both the initial stimulation geometry and the subsequent diffusion of feedback molecules to neighboring unstimulated cells. By iteration 500, the spatial pattern exhibits clear concentration gradients extending from the initially stimulated regions, demonstrating the platform’s ability to capture both direct cellular responses and indirect paracrine effects.
@@ Discussion
ECMSim advances computational systems biology through its integration of high-performance numerical simulation with accessible web-based deployment, addressing fundamental limitations that have historically restricted the adoption of sophisticated modeling platforms in cardiovascular research. The platform’s implementation of over 1 million coupled ordinary differential equations within a browser environment demonstrates the maturation of WebAssembly technology for scientific computing applications, achieving computational performance within 20% of native implementations while eliminating the substantial barriers associated with specialized software installation and configuration. The mathematical framework underlying ECMSim advances beyond traditional approaches through its explicit modeling of spatial heterogeneity in cardiac fibroblast populations, incorporating both intracellular signaling complexity and intercellular communication through reaction-diffusion dynamics. This multi-scale integration addresses a critical gap identified in recent reviews of cardiac fibrosis modeling, 23, 13, 25 where spatial organization and cell-cell communication have been largely overlooked despite their central importance in pathological remodeling processes. The platform’s brush-based spatial input system enables precise experimental design that mirrors clinical intervention strategies, such as localized drug delivery or regional mechanical stimulation protocols. 26, 1 The implementation of multiplicative regulatory interactions rather than simplified linear kinetics represents a significant advancement in capturing the complex nonlinear behaviors that characterize cellular signaling networks. This approach aligns with recent experimental findings demonstrating the prevalence of cooperative and competitive interactions in fibroblast signaling pathways, 27, 28 providing a more biologically realistic foundation for understanding emergent network behaviors and therapeutic intervention strategies.
@@
The simulation results presented in Figure 4 reveal fundamental principles of spatiotemporal pattern formation in cardiac tissue that have important implications for understanding pathological progression. The emergence of concentration gradients extending beyond initially stimulated regions demonstrates how localized inflammatory or mechanical stimuli can propagate across tissue domains through paracrine signaling mechanisms, consistent with recent experimental observations of fibrotic spread following myocardial infarction. 29, 30 The platform’s ability to capture both direct cellular responses and indirect paracrine effects addresses a fundamental challenge in translating molecular-scale experimental findings to tissue-level pathophysiology. Recent single-cell RNA sequencing studies have revealed substantial heterogeneity in cardiac fibroblast populations, 31, 32 and ECMSim’s spatial framework provides a computational environment for investigating how this heterogeneity contributes to tissue-level function and dysfunction.
@@
ECMSim’s real-time parameter adjustment capabilities enable systematic investigation of therapeutic intervention strategies through virtual drug screening approaches. The platform’s ability to simulate localized drug application through the spatial brush interface directly parallels emerging clinical approaches such as targeted nanoparticle delivery systems and localized gene therapy protocols 33. The comprehensive pathway coverage enables investigation of both on-target and off-target effects, addressing a critical limitation in current drug development pipelines where systemic effects are often poorly predicted from single-pathway studies. 34 The platform’s capacity to model combinatorial interventions provides particular value for developing multi-target therapeutic strategies that may be necessary for effective fibrosis treatment, suggesting that successful interventions will require simultaneous modulation of multiple pathways. ECMSim enables systematic exploration of combination therapy effects while accounting for potential antagonistic interactions that could limit therapeutic efficacy. The integration of mechanotransduction pathways enables investigation of mechanical intervention strategies, including cardiac rehabilitation protocols and device-based therapies. The platform can model how changes in mechanical loading conditions influence cellular signaling and matrix remodeling, providing insights into optimal timing and intensity of mechanical interventions. 35
@@
The WebAssembly-based architecture represents a significant technological advancement that addresses longstanding barriers to computational tool adoption in biological research. Traditional modeling platforms require specialized software installation, computational expertise, and often substantial computational resources that limit accessibility to researchers with advanced technical backgrounds. ECMSim’s browser-based deployment eliminates these barriers while maintaining computational rigor, potentially democratizing access to sophisticated modeling capabilities across diverse research environments. The platform’s modular architecture facilitates community-driven development through its object-oriented design and open-source implementation. This approach aligns with recent calls for increased reproducibility and transparency in computational biology, 36 enabling researchers to examine, modify, and extend the underlying mathematical models. The web-based deployment facilitates rapid dissemination of model updates and community contributions, addressing the version control challenges that often plague collaborative model development efforts. The real-time visualization capabilities provide immediate feedback that transforms the traditional modeling workflow from a batch-processing approach to an interactive exploration paradigm. This shift has important implications for educational applications, enabling students and trainees to develop intuitive understanding of complex signaling networks through direct manipulation and observation.
@@
Several aspects of the current implementation warrant consideration for future enhancement. The two-dimensional spatial grid, while computationally efficient, represents a simplification of the complex three-dimensional architecture characteristic of cardiac tissue. Extension to three-dimensional geometries would enable investigation of transmural signaling gradients and the influence of tissue geometry on pattern formation, phenomena that are increasingly recognized as important factors in cardiac remodeling. 23 The current implementation assumes homogeneous fibroblast populations, neglecting the phenotypic diversity revealed by recent single-cell studies. 31, 37 Integration of cell-type-specific parameter sets based on single-cell transcriptomic data would enhance biological realism and enable investigation of how cellular heterogeneity contributes to tissue-level responses. This enhancement could incorporate stochastic elements to capture the noise inherent in cellular signaling systems, particularly relevant for low-abundance molecular species where fluctuations may significantly influence system behavior. The deterministic nature of the current mathematical framework may not adequately capture the stochastic fluctuations that characterize cellular signaling networks. Integration of stochastic differential equation approaches or agent-based modeling components could provide more realistic representations of cellular noise and its effects on pattern formation and signal propagation. 38 Future developments could incorporate patient-specific parameterization through integration with clinical data sources, enabling personalized modeling approaches that account for individual genetic backgrounds, comorbidity profiles, and treatment histories. This capability would support precision medicine approaches by predicting individual responses to specific therapeutic interventions. 39
@@
The technological and methodological innovations demonstrated in ECMSim have implications extending beyond cardiac fibrosis to diverse areas of biomedical research. The reaction-diffusion framework underlying the platform is fundamentally applicable to any biological system involving spatial pattern formation and intercellular communication, including cancer progression, wound healing, developmental morphogenesis, and tissue engineering applications. 40 The platform’s success in democratizing access to sophisticated computational tools suggests potential for broader transformation of computational biology education and research practices. The combination of mathematical rigor with intuitive interfaces could accelerate the integration of quantitative approaches into traditionally experimental research programs, fostering interdisciplinary collaboration and accelerating scientific discovery. ECMSim’s demonstration that complex biological systems can be effectively modeled within standard web browsers opens new possibilities for collaborative research across institutional boundaries. The platform could serve as a foundation for developing specialized modules targeting specific disease processes or therapeutic interventions, creating an ecosystem of interconnected modeling tools that collectively address the complexity of human disease. 41 The open-source architecture and community-driven development model position ECMSim to evolve continuously in response to advancing biological knowledge and technological capabilities. This approach ensures long-term sustainability while enabling rapid incorporation of emerging experimental findings into the computational framework, maintaining relevance as the field advances.
@@
The original code is available at Richardson Lab in GitHub (https://github.com/SysMechBioLab/ECMSim)

### v2

@@ INTRODUCTION
Screenshots of 3 distinct cases of simulation. All the matrices have collagen I in the ECM and simulated 4 different iterations (0, 100, 300, 500) with 0.1 time steps (S0, simulation times 0, 10, 30, and 50 respectively). More yellow color represents the higher concentration and black color represents the lowest concentration in the matrix. Case 01: No brush selected. So, the inputs molecules have default values (all at 0). Case 02: Brush selected (X shape). TGF- input value is 1 in the brushed area. Rest of the inputs are in default values (0). Unselected area has default input molecule concentration. Case 03: Brush selected (X shape). TGF-, AngII, IL6, IL1, TNF, and NE input molecules are set at 1. Unselected area has default input molecule concentration. Each brush selected cases (2 and 3 - each case separately) included same brush selection throughout the continuous temporal iterations.
@@ RESULTS
ECMSim successfully implements a comprehensive computational framework for simulating cardiac ECM remodeling through the integration of 137 molecular species and 259 regulatory interactions distributed across a spatially-resolved tissue grid. Fig. 2A illustrates the fundamental feedback loop architecture underlying the model, where external TGF- input triggers intracellular signaling cascades that ultimately produce both ECM molecules and feedback molecules that diffuse to neighboring cells. This autocrine and the paracrine signaling mechanism enables the propagation and amplification of fibrotic responses through the tissue. The complete signaling network architecture (Fig. 2B) encompasses the major pathways implicated in cardiac fibrosis, including TGF-/Smad signaling, MAPK cascades, PI3K-Akt-mTOR pathways, and mechanotransduction networks. The hierarchical organization of the network enables signal integration from multiple upstream inputs through intermediate processing nodes to downstream effector molecules responsible for matrix protein synthesis. The modular architecture facilitates systematic investigation of pathway crosstalk and the identification of critical regulatory nodes that control fibrotic responses.
The spatial implementation (Fig. 2C) demonstrates the platform’s capability to model tissue-level organization through a regular cellular grid where each cell contains the complete intracellular signaling network. Intercellular communication is mediated through the diffusion of five key feedback molecules that coordinate regional responses and enable the emergence of spatially heterogeneous patterns. The diffusion dynamics (Fig. 2D) show representative simulation outputs displaying the spatial distribution of molecular concentrations at successive time points, illustrating how localized stimuli can propagate across the tissue domain through diffusive coupling. To further characterize the spatial diffusion behavior, Fig. 2E–G presents vector field streamlines, gradient fields, and three-dimensional surface visualizations for four discrete concentration sources placed at different grid locations. The vector field (Fig. 2E) illustrates the direction and magnitude of diffusive flux, while the gradient field (Fig. 2F) and 3D surface (Fig. 2G) show how multiple diffusion sources interact to form composite concentration landscapes with smooth spatial gradients.
In ECMSim, the fibrotic response can be quantified through multiple complementary metrics: (1) collagen accumulation, specifically the concentrations of procollagen type I (proCI) and procollagen type III (proCIII), which are the primary structural ECM proteins in cardiac fibrosis; (2) the collagen/MMP balance, representing the ratio of collagen production to MMP-mediated degradation; (3) myofibroblast activation markers, particularly -SMA expression; and (4) overall ECM protein accumulation across all 17 modeled species. While collagen accumulation is the most commonly used single metric for fibrosis severity, the platform enables multi-dimensional characterization of the fibrotic response, allowing researchers to select the most appropriate output variables for their specific research questions.
@@
The ECMSim web application provides an intuitive interface for exploring complex signaling dynamics through multiple integrated visualization modalities (Fig. 3). The primary heatmap visualization displays the spatial distribution of selected molecular species across the 100×100 cellular grid, with real-time updates during simulation execution that enable immediate observation of evolving concentration patterns. The molecule selection panel allows users to dynamically switch between 17 ECM molecules and 4 feedback molecules, instantly updating both spatial visualizations and temporal plots without requiring simulation restart. The innovative brush selection tool enables precise spatial control over input conditions by allowing users to define arbitrary cell populations for targeted interventions. The brush interface supports adjustable sizing from single-cell precision to broad regional selections, accommodating diverse experimental scenarios ranging from localized growth factor application to regional mechanical stimulation. The input concentration control panel provides independent adjustment of 10 primary signaling molecules, with real-time value displays and immediate propagation to selected cell populations. This design enables complex experimental scenarios such as spatially heterogeneous inflammation conditions or combinatorial cytokine treatments.
The visualization system combines tissue-level spatial patterns with detailed single-cell temporal dynamics. Users can select up to eight individual cells directly on the main heatmap by clicking, with their molecular trajectories displayed in real-time line plots that enable direct comparison between cells experiencing different microenvironments. Selected cells can be individually edited, allowing users to manually adjust the concentration of any molecular species in a specific cell to test localized perturbations. The platform also provides figure export functionality, enabling users to save the current heatmap visualization in PNG and SVG formats for publication or further analysis. The comprehensive parameter control panel contains all eight primary rate constants with real-time adjustment capabilities, enabling immediate exploration of model sensitivity and behavior under different kinetic assumptions. This interactive capability transforms ECMSim into a digital laboratory environment where hypotheses can be tested and validated without the temporal and resource constraints of traditional experimental approaches.
@@
The platform’s spatial modeling capabilities enable investigation of emergent tissue-level phenomena through the simulation of diffusion-driven pattern formation. Fig. 4 demonstrates three representative simulation scenarios illustrate the platform’s ability to model spatially heterogeneous responses to localized stimuli. Case 01 represents the baseline condition with uniform zero input concentrations across all cells, resulting in spatially homogeneous concentrations dominated by stochastic fluctuations around the initial random values. The temporal evolution from iteration 0 to 500 shows gradual equilibration toward steady-state values with minimal spatial organization.
Case 02 demonstrates the response to localized TGF- stimulation applied in an X-shaped pattern through the brush selection interface. The TGF- input concentration of 1.0 in the selected region triggers robust intracellular signaling cascades that drive procollagen I (proCI) synthesis in the stimulated cells. During the early transient phase (e.g., iteration 100), the stimulated X region may display lower proCI concentration than the surrounding unstimulated cells because the intracellular signaling cascades are actively transitioning from the randomized initial state through receptor activation and kinase processing before ECM production increases; this lag reflects the time required for signals to propagate through the multi-step pathway hierarchy. At later time points (iterations 300–500), the stimulated region shows markedly higher concentrations as the signaling cascades have fully propagated to ECM production outputs. The temporal evolution reveals the gradual emergence of spatial patterns that reflect both the initial stimulation geometry and the subsequent diffusion of feedback molecules to neighboring unstimulated cells. By iteration 500, the spatial pattern exhibits clear concentration gradients extending from the initially stimulated regions. With continuous brush-applied stimulation, the system reaches a spatially heterogeneous steady state that maintains these gradients due to the persistent localized source; without sustained stimulation, the system would eventually approach a uniform steady state as diffusion equilibrates the concentration field, though this requires simulation times longer than the diffusion timescale across the full grid.
@@
Parameter sensitivity analysis
To identify the dominant regulators of ECM remodeling outcomes, a one-at-a-time (OAT) sensitivity analysis was performed on the eight primary adjustable rate constants (, , , , , , , and the integration time step ). Each parameter was perturbed by and from its default value while all other parameters were held constant. The sensitivity index was computed as the relative change in the transient area under the curve (AUC) of each ECM output during the dynamic accumulation phase (–), where parameter differences manifest most clearly. Seven key ECM outputs were tracked: procollagen I, procollagen III, MMP-2, MMP-9, TIMP-1, -SMA, and fibronectin.
Fig. 5A presents the sensitivity heatmap, where each column corresponds to a perturbation level of a given parameter and each row represents an ECM output. The analysis reveals that the ECM production rate () is by far the most influential parameter (mean ), followed by the activation rate (, mean ) and the degradation rate (, mean ). The tornado chart (Fig. 5B) confirms this hierarchy for procollagen I, showing that perturbation of produces sensitivity indices of approximately , while all other parameters produce indices below . The remaining parameters (, , , , and ) have comparatively minor effects on ECM outputs, indicating that the model is robust to moderate variations in these parameters. These findings are consistent with the biological understanding that ECM accumulation in cardiac fibrosis is primarily governed by the rate of matrix protein production by activated fibroblasts and the balance between activation and degradation kinetics.
@@ DISCUSSION
ECMSim advances computational systems biology through its integration of high-performance numerical simulation with accessible web-based deployment, addressing fundamental limitations that have historically restricted the adoption of sophisticated modeling platforms in cardiovascular research. The platform’s implementation of over 1 million coupled ordinary differential equations within a browser environment demonstrates the maturation of WebAssembly technology for scientific computing applications, achieving computational performance within 20% of native implementations while eliminating the substantial barriers associated with specialized software installation and configuration. The mathematical framework underlying ECMSim advances beyond traditional approaches through its explicit modeling of spatial heterogeneity in cardiac fibroblast populations, incorporating both intracellular signaling complexity and intercellular communication through reaction-diffusion dynamics. This multi-scale integration addresses a critical gap identified in recent reviews of cardiac fibrosis modeling, 45, 46, 33, 34, 32, 43, 14 where spatial organization and cell-cell communication have been largely overlooked despite their central importance in pathological remodeling processes. The platform’s brush-based spatial input system enables precise experimental design that mirrors clinical intervention strategies, such as localized drug delivery or regional mechanical stimulation protocols. 9, 41 The implementation of multiplicative regulatory interactions rather than simplified linear kinetics represents a significant advancement in capturing the complex nonlinear behaviors that characterize cellular signaling networks. This approach aligns with recent experimental findings demonstrating the prevalence of cooperative and competitive interactions in fibroblast signaling pathways, 40, 11 providing a more biologically realistic foundation for understanding emergent network behaviors and therapeutic intervention strategies.
@@
The simulation results presented in Fig. 4 reveal fundamental principles of spatiotemporal pattern formation in cardiac tissue that have important implications for understanding pathological progression. The emergence of concentration gradients extending beyond initially stimulated regions demonstrates how localized inflammatory or mechanical stimuli can propagate across tissue domains through paracrine signaling mechanisms, consistent with recent experimental observations of fibrotic spread following myocardial infarction. 23, 10 The platform’s ability to capture both direct cellular responses and indirect paracrine effects addresses a fundamental challenge in translating molecular-scale experimental findings to tissue-level pathophysiology. Recent single-cell RNA sequencing studies have revealed substantial heterogeneity in cardiac fibroblast populations, 37, 28 and ECMSim’s spatial framework provides a computational environment for investigating how this heterogeneity contributes to tissue-level function and dysfunction.
@@
ECMSim’s real-time parameter adjustment capabilities enable systematic investigation of therapeutic intervention strategies through virtual drug screening approaches. The platform’s ability to simulate localized drug application through the spatial brush interface directly parallels emerging clinical approaches such as targeted nanoparticle delivery systems and localized gene therapy protocols 18. The comprehensive pathway coverage enables investigation of both on-target and off-target effects, addressing a critical limitation in current drug development pipelines where systemic effects are often poorly predicted from single-pathway studies. 48 The platform’s capacity to model combinatorial interventions provides particular value for developing multi-target therapeutic strategies that may be necessary for effective fibrosis treatment, suggesting that successful interventions will require simultaneous modulation of multiple pathways. ECMSim enables systematic exploration of combination therapy effects while accounting for potential antagonistic interactions that could limit therapeutic efficacy. The integration of mechanotransduction pathways enables investigation of mechanical intervention strategies, including cardiac rehabilitation protocols and device-based therapies. The platform can model how changes in mechanical loading conditions influence cellular signaling and matrix remodeling, providing insights into optimal timing and intensity of mechanical interventions. 6
@@
The WebAssembly-based architecture represents a significant technological advancement that addresses longstanding barriers to computational tool adoption in biological research. Traditional modeling platforms require specialized software installation, computational expertise, and often substantial computational resources that limit accessibility to researchers with advanced technical backgrounds. ECMSim’s browser-based deployment eliminates these barriers while maintaining computational rigor, potentially democratizing access to sophisticated modeling capabilities across diverse research environments. The platform’s modular architecture facilitates community-driven development through its object-oriented design and open-source implementation. This approach aligns with recent calls for increased reproducibility and transparency in computational biology, 38 enabling researchers to examine, modify, and extend the underlying mathematical models. The web-based deployment facilitates rapid dissemination of model updates and community contributions, addressing the version control challenges that often plague collaborative model development efforts. The real-time visualization capabilities provide immediate feedback that transforms the traditional modeling workflow from a batch-processing approach to an interactive exploration paradigm. This shift has important implications for educational applications, enabling students and trainees to develop intuitive understanding of complex signaling networks through direct manipulation and observation.
@@
Several aspects of the current implementation warrant consideration for future enhancement. The “tension” variable in the current model represents mechanical stimulation that induces molecular- scale force across cell adhesion proteins and ion channels. Fibroblast mechanotransduction can driven by extracellular deformation and/or intracellular acto-myosin contractility, which activate downstream signaling cascades and ultimately influence gene expression through transcription factors such as YAP/TAZ and SRF. The “tensionfb” feedback variable provides a way for our model to capture a sort of force propagation where tension from one cell pulling on their local matrix can act on other cells within small distances. Note we are not actually solving detailed traction force balances. A full finite element mechanical model coupled with the biochemical signaling network would be a valuable future extension that would enable modeling of physical deformations, fiber reorientation, and direct force transmission through the ECM. The current implementation supports periodic boundary conditions as the default, which effectively approximate an infinite tissue domain. Alternative boundary conditions, including fixed-value (Dirichlet) and no-flux (Neumann) conditions, would be valuable for modeling specific physiological scenarios such as tissue boundaries and infarct borders, and the modular code architecture readily supports their addition in future releases.
The two-dimensional spatial grid, while computationally efficient, represents a simplification of the complex three-dimensional architecture characteristic of cardiac tissue. Extension to three-dimensional geometries would enable investigation of transmural signaling gradients and the influence of tissue geometry on pattern formation, phenomena that are increasingly recognized as important factors in cardiac remodeling. 45 The current implementation assumes homogeneous fibroblast populations, neglecting the phenotypic diversity revealed by recent single-cell studies. 37, 8 Integration of cell-type-specific parameter sets based on single-cell transcriptomic data would enhance biological realism and enable investigation of how cellular heterogeneity contributes to tissue-level responses. This enhancement could incorporate stochastic elements to capture the noise inherent in cellular signaling systems, particularly relevant for low-abundance molecular species where fluctuations may significantly influence system behavior. The deterministic nature of the current mathematical framework may not adequately capture the stochastic fluctuations that characterize cellular signaling networks. Integration of stochastic differential equation approaches or agent-based modeling components could provide more realistic representations of cellular noise and its effects on pattern formation and signal propagation. 24 Future developments could incorporate patient-specific parameterization through integration with clinical data sources, enabling personalized modeling approaches that account for individual genetic backgrounds, comorbidity profiles, and treatment histories. This capability would support precision medicine approaches by predicting individual responses to specific therapeutic interventions. 1
@@
The technological and methodological innovations demonstrated in ECMSim have implications extending beyond cardiac fibrosis to diverse areas of biomedical research. The reaction-diffusion framework underlying the platform is fundamentally applicable to any biological system involving spatial pattern formation and intercellular communication, including cancer progression, wound healing, developmental morphogenesis, tissue-based transcriptomics, 13 and tissue engineering applications. 26 The platform’s success in democratizing access to sophisticated computational tools suggests potential for broader transformation of computational biology education and research practices. The combination of mathematical rigor with intuitive interfaces could accelerate the integration of quantitative approaches into traditionally experimental research programs, fostering interdisciplinary collaboration and accelerating scientific discovery. ECMSim’s demonstration that complex biological systems can be effectively modeled within standard web browsers opens new possibilities for collaborative research across institutional boundaries. The platform could serve as a foundation for developing specialized modules targeting specific disease processes or therapeutic interventions, creating an ecosystem of interconnected modeling tools that collectively address the complexity of human disease. 16 The open-source architecture and community-driven development model position ECMSim to evolve continuously in response to advancing biological knowledge and technological capabilities. This approach ensures long-term sustainability while enabling rapid incorporation of emerging experimental findings into the computational framework, maintaining relevance as the field advances.
@@
The original code is available at Richardson Lab in GitHub (https://github.com/SysMechBioLab/ECMSim).